jimmy walks from school 1.5km west in 0.50h to his grandmothers for 1.0h. he then walk home east 2.5km in 1.5h explain answer
A. what is his total displacement
B. what is his total distance traveled
C. what is his total time
D. what is jimmy's average speed
E. what is Jimmy's average velocity
1
Expert's answer
2020-11-10T09:59:19-0500

Assume that when he walks east, it is positive displacement, when he wals west, this is negative displacement:

A. His total displacement is


"A=d_W+d_E=-1.5+2.5=1\text{ km}."

B. His total distance traveled is


"B=|d_W|+|d_E|=1.5+2.5=4\text{ km}."

C. His total time is


"t_C=1+1.5=2.5\text{ h}."

D. Jimmy's average speed is


"s=\frac{B}{t_C}=1.6\text{ km/h}."

E. Jimmy's average velocity is


"v=\frac{A}{t_c}=0.4\text{ km/h}."